Boxes are made by cutting 8 cm squares from the corners of sheets of cardboard and then folding. The sheets of cardboard are 6 cm longer than they are wide.
For the sheet of cardboard whose width is x cm, find expressions, in terms of xm for:
(i)(a) Length of the sheet
b) The length of the finished box
c) Width of the finished box
(ii)Show that the volume of the box is 8x^2 - 208x + 1280cm^3

Answer by mananth(16946)

let width of sheet be x cm
length = x+6cm
...
width of box = x-16 cm
width of sheet = x+6-16=x-10cm
height = 8
....
Volume of box = L*W*H
(x-16)(x-10)*8
(x^2-26x+160)*8
8x^2-208x+1280 cm^3